Consider the following quadratic equation: y = 5x 2 + 15x − 50.


What are the coordinates of the point(s) where the graph of the function crosses the x-axis?


(a) (0, −2) and (0, 4)


(b) (0, 2) and (0, 4)


(c) (2, 0) and (−5, 0)


(d) (2, 0) and (4, 0)


1
Expert's answer
2021-11-15T14:29:05-0500
"y = 5x^ 2 + 15x \u2212 50"

"y=0=> 5x^ 2 + 15x \u2212 50=0"

"x^ 2 + 3x \u2212 10=0"

"(x+5)(x-2)=0"

"x_1=-5, x_2=2"

(c) (2, 0) and (−5, 0)